Dublin University Rifle Club

Version 1.1 Date : May 4, 2001 Filename : /home/dennehym/DURC/Documents/Inventory/RCS/Inventory2001.tex,v Status : Exp

Dublin University Rifle Club Equipment Inventory

Mark Dennehy

May 4, 2001

1 Cleaning Supplies

1× M-74 Ammunition box, for storage